Compute the yield to maturty the new issue. ftertax cost of debt tstanding with grow a. b. te tax adjustment to determine the aftertax co Make the appropria 30 years to maturity. The bond carries an annual interest payment currently selling for $880 per bond. Russell Corp. is in a 40 percer The firm wishes to know what the aftertax cost of a new bond is of $105 and ed t and cost of Russell Container Corporation has a $1.000 par value bai ely to ield to matu- riy on the old issue because the risk and maturity date will be sin a Compute the yield to maturity on the old issue and use this as be. The yield to maturity on the new issue will be the same as the y for the new issue Make the appropriate tax adjustment to determine the aftertax c cost of debt. areent tax bracket and has a bond outstanding
Answer #1

a. We have following formula for calculation of bond’s yield to maturity (YTM)

Bond price P0 = C* [1- 1/ (1+YTM) ^n] /i + M / (1+YTM) ^n

Where,

P0 = the current market price of bond = $880 per bond

C = coupon payment = $105

n = number of payments (years to maturity) = 30

YTM = interest rate, or yield to maturity =?

M = value at maturity, or par value = $ 1000

Now we have,

$ 880 = $105 * [1 – 1 / (1+YTM) ^30] /i + 1000 / (1+YTM) ^30

By trial and error method we can calculate the value of YTM = 11.99% per year

[Or you can use excel function for YTM calculation in following manner

“= Rate(N,PMT,PV,FV)”

“Rate(30,-105,880,-1000)” = 11.99%]

b. The before tax cost of debt is the Yield to maturity (YTM) of the bond which is 11.99%

The after tax cost of debt = YTM *(1 – T)

Where T is the tax rate =40%

And YTM = 11.99%

Therefore

After tax cost of debt = 11.99% *(1 – 0.40) = 7.193%